OpenStudy (anonymous):

The equation 14m = 400.40 can be used to find the cost per yard m, in dollars, of mulch. What is the value of m?

sammixboo (sammixboo):

All you need to do is 400.40/14

OpenStudy (anonymous):

100.1?

sammixboo (sammixboo):

Here.. You need to make m by itself. To do do divide 14 on both sides. 14/14m = 400.40/14 m = 400.40/14

sammixboo (sammixboo):

Nope! Try again

sammixboo (sammixboo):

14 x 100.1 = 1,401.4

sammixboo (sammixboo):

So that's not it

OpenStudy (anonymous):

oohh

sammixboo (sammixboo):

So did you get 400.40/14?

OpenStudy (anonymous):

no still trying

sammixboo (sammixboo):

Okay take your time! No one else needs my help right now :)

OpenStudy (anonymous):

sigh i keep getting 100.1

sammixboo (sammixboo):

Hmm..

sammixboo (sammixboo):

Are you sure you are putting 400 . 40 / 14

OpenStudy (anonymous):

dividing right ?

sammixboo (sammixboo):

Yeppers

OpenStudy (anonymous):

28.6

sammixboo (sammixboo):

Correct! m = 28.6!

sammixboo (sammixboo):

14m = 400.40 14 * 28.6 = 400.40 400.40 = 400.40

OpenStudy (anonymous):

awesome

sammixboo (sammixboo):

Yeah =)

sammixboo (sammixboo):

If you have anymore question tag me \(\rm \color{red}{@}\color{magenta}{s}\color{red}{a}\color{magenta}{m}\color{red}i\color{magenta}x\color{red}b\color{magenta}{o}\color{red}o\)

sammixboo (sammixboo):

Omg I spelled my name wrong

sammixboo (sammixboo):

LOL @arabpride

sammixboo (sammixboo):

I mean @sammixboo not @samixboo

OpenStudy (shadowlegendx):

@sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo @sammixboo

OpenStudy (shadowlegendx):

Put one of those^

sammixboo (sammixboo):

o_o

OpenStudy (slinley):

You're amazing and smart ^_^ @sammixboo